OpenStudy (anonymous):

Sgadi, can you show where to plug in the point and steps how to solve this equation? Slope Intercept Form of the equation of a line that passes the point (3,-2) and had a slope of 1/3. Formula which you can use is y=mx+c where "m" is slope then put the point in the equation and get the value of c. c=y−mx where (x,y) is given point.

OpenStudy (sgadi):

using formula \[y=mx+(q−mp)\] here m=1/3, p=3, q=-2 \[y=frac{1}{3}x+(-2-3\frac{1}{3})\] \[y=frac{1}{3}x-3\] \[3y=x-9\] \[x-3y-9=0\]

OpenStudy (anonymous):

Sgadi, thank you very much for the assistance

OpenStudy (sgadi):

welcome